# EIKE Workshop at JOWO

### v1

@@ 4 Implications for computational metaphor processing and future work
Datasets. Metaphor datasets should not only account for simple domain mappings within sentences, but also adopt the possibility to annotate conversational data accounting for dimensions other than the cognitive one. Furthermore, datasets concerning perlocutionary effects should also link to ontological resources about emotions. Additionally, no dataset yet accounts for the fact that meaning can be represented as a spectrum, thus future work should account for this conception of meaning. Finally, employing diverse annotators is fundamental in the view employed in this paper.
@@
Knowledge representation. Current ontological frameworks for metaphor representation, as described in Section 2, have made important strides by encoding source–target mappings and blend spaces. However, they tend to treat metaphor as a phenomenon abstracted from discourse and pragmatic intent. In contrast, our proposal emphasizes a multi-layered model of meaning that integrates literal, conceptual, and pragmatic levels, requiring a representational shift that can, in future developments of this work, be ontologically formalized. In fact, metaphor representation should move toward stratified ontological models that includes a pragmatic layer capturing the illocutionary and perlocutionary acts of metaphorical utterances, grounded in discourse context, communicative intent, and emotional effects.
Finally, such a layered representation can support updates or reinterpretations of metaphorical meaning in conversational settings or evolving discourse. This necessitates reasoning mechanisms capable of tracking metaphor reinterpretation and resistance, grounded in community norms and individual differences.
@@
Computational metaphor processing. The proposed framework has direct implications for computational processing. By aligning processing tasks with the three dimensions, we can design systems that generate, interpret, and apply metaphors in a more context-aware, human-aligned, and socially beneficial manner. For instance, as a first step, both the LAG approach [25] and the METCL [35] approach can be employed in a complementary way to assess the conceptual combination layer for metaphor processing. We argue that METCL and LAG are not only compatible but potentially integrable and mutually reinforcing. Specifically, LAG could leverage the prototypical property lists provided by METCL to weight RDF triples, yielding more precise definitions and reducing LLM hallucinations. Conversely, METCL could exploit LAG’s capacity to identify missing frames in MetaNet or to align typical features with ontological categories, thus enabling a shared RDF layer for interoperability. For instance, LAG might detect objects in an image, while METCL could suggest which visual attributes instantiate their prototypical properties.
@@
Metaphor understanding, too, can benefit from the layered representation we propose. By encoding both conceptual mappings and pragmatic implications, systems can disambiguate metaphorical usage more accurately, especially in real-life settings.
LLMs, when coupled with structured semantic representations, may serve as effective metaphor interpreters. In such a neurosymbolic setup, the literal surface form is parsed, mapped to conceptual domains, and then interpreted through pragmatic filters reflecting the speaker’s goal, emotional tone, and discourse situation. This facilitates better performance in metaphor detection, paraphrasing, and explanation tasks, especially in underexplored genres like dialogue, narratives, or scientific texts.
Metaphors for social good. Metaphors are not neutral; they carry social, cultural and emotional weight, and using or detecting them responsibly has implications across several socially relevant domains (see Section 2). In education, adaptive, metaphor-aware tools could tailor figurative language to a learner’s background knowledge, helping students grasp abstract ideas more easily. For science communication, apt metaphors make complex phenomena more engaging and accessible to the public by matching scientific concepts with images that resonate pragmatically and emotionally. In doctor-patient interactions, the metaphors chosen—such as describing cancer as an invader versus a malfunction—can reshape trust, empathy and the overall therapeutic experience. Legal and policy discourse is likewise shaped by metaphor, so computational systems that identify and reframe figurative language in statutes or briefs can promote clearer and fairer interpretations. Finally, hate speech frequently relies on dehumanizing metaphors; recognizing and defusing such figurative frames (including in multimodal content like memes) is essential for effective moderation and the mitigation of social harm. Our proposed framework acknowledges the ethical consequences metaphors have in real-life settings.
@@
5 Conclusion
Metaphors are cognitive tools deeply embedded in human reasoning and communication. In this work, we assess current problems in metaphor processing and propose a three-layered framework for metaphor processing, encompassing literal content, conceptual blending, and pragmatic intent. This onion-like stratification allows computational models to capture not only what metaphors mean but what they do—illuminating both their cognitive structure and their communicative function. Although the operational framework is currently theoretical, we show how it can serve both as a fruitful direction for future research and as a foundation for subsequent empirical work on this issue.
Our approach bridges key theoretical traditions—from Conceptual Blending Theory to Speech Act Theory—while offering practical implications for the design of computational systems that are metaphor-aware, context-sensitive, and socially responsible. By formalizing metaphor as a structured, multi-dimensional phenomenon, we lay the foundation for new datasets, evaluation metrics, and neurosymbolic methods that better reflect the richness of metaphor in natural discourse.
Looking ahead, we envision extending this framework to new modalities and domains and implementing it in computational experiments.

### v2

@@ 4 Implications for computational metaphor processing and future work
Datasets. Metaphor datasets should not only account for simple domain mappings within sentences, but also adopt the possibility to annotate conversational and multimodal data accounting for dimensions other than the cognitive one. To date, no single metaphor-related dataset annotates attitude, speech acts and clues of utterances, but initial effort is being made towards this step, for example in the representation of intentions [43]. Furthermore, datasets concerning perlocutionary effects should also link to ontological resources about emotions. Additionally, no dataset yet accounts for the fact that meaning can be represented as a spectrum, thus future work should account for this conception of meaning. Finally, employing diverse annotators is fundamental in the view employed in this paper, in order to obtain the first gold standard fine-grained pragmatic datasets in the field.
@@
Knowledge representation. Current ontological frameworks for metaphor representation, as described in Section 2, have made important strides by encoding source–target mappings and blend spaces. However, they tend to treat metaphor as a phenomenon abstracted from discourse and pragmatic intent. In contrast, our proposal emphasizes a multi-layered model of meaning that integrates literal, conceptual, and pragmatic levels, requiring a representational shift that can, in future developments of this work, be ontologically formalized. In fact, metaphor representation should move toward stratified ontological models that includes a pragmatic layer capturing the illocutionary and perlocutionary acts of metaphorical utterances, grounded in discourse context, communicative intent, and emotional effects. Finally, such a layered representation can support updates or reinterpretations of metaphorical meaning in conversational settings; for instance, through reasoning mechanisms capable of tracking metaphor reinterpretation and resistance, grounded in community norms and individual differences.
@@
Possible application of the framework in computational metaphor processing. The proposed framework has direct implications for computational processing. By aligning processing tasks with the three dimensions, we can design systems that generate, interpret, and apply metaphors in a more context-aware, human-aligned, and socially beneficial manner. For instance, both the LAG approach [25] and the METCL [35] approach can be employed in a complementary way to assess the conceptual content and combination layer for metaphor processing. We argue that METCL and LAG are not only compatible but potentially integrable and mutually reinforcing. Specifically, LAG could leverage the prototypical property lists provided by METCL to weight RDF triples, yielding more precise definitions and reducing LLM hallucinations. Conversely, METCL could exploit LAG’s capacity to identify missing frames in MetaNet or to align typical features with ontological categories, thus enabling a shared RDF layer for interoperability. For instance, LAG might detect objects in an image, while METCL could suggest which visual attributes instantiate their prototypical properties. For what concerns the pragmatic layer, the lack of datasets currently hinders computational implementation (see Section4), starting from a gold-standard annotated dataset, but LLM-based approaches such as LAG can be used to extract pragmatic features, which can in turn be analyzed.
@@
Metaphor understanding, too, can benefit from the layered representation we propose. By encoding both conceptual mappings and pragmatic implications, systems can disambiguate metaphorical usage more accurately, especially in real-life settings. LLMs, when coupled with structured semantic representations, may serve as effective metaphor interpreters. In such a neurosymbolic setup, the literal surface form is parsed, mapped to conceptual domains, and then interpreted through pragmatic filters reflecting the speaker’s goal, emotional tone, and discourse situation. This facilitates better performance in metaphor detection, paraphrasing, and explanation tasks, especially in underexplored genres like dialogue, narratives, or scientific texts.
Metaphors for social good. Metaphors are not neutral; they carry social, cultural and emotional weight, and using or detecting them responsibly has implications across several socially relevant domains (see Section 2), for instance employing metaphor-aware computational tools in education, science communication, doctor-patient interactions and hate speech.
@@
5 Limitations and conclusion
Metaphors are cognitive tools deeply embedded in human reasoning and communication. In this work, we assess current problems in metaphor processing and propose a three-layered framework for metaphor processing, encompassing conceptual and pragmatic implicit knowledge together. This onion-like stratification allows computational models to capture not only what metaphors mean but what they do, illuminating both their cognitive structure and their communicative function. Although the operational framework is currently theoretical, we show how it can serve both as a fruitful direction for future research and as a foundation for subsequent empirical work on this issue, for instance, for the design of computational systems that are metaphor-aware, context-sensitive, and socially responsible. By formalizing metaphor as a structured, multi-dimensional phenomenon, we lay the foundation for new datasets, evaluation metrics, and neurosymbolic methods that better reflect the richness of metaphor in natural discourse. The main current limitation of this framework is the lack of computational implementation. It is also necessary to develop a robust and shared line of research that integrates CBT and pragmatics theories on metaphor interpretation, accompanied by the discussion of a common vocabulary for the construction of future datasets, and, on the computational side, by a coherent formalization of pragmatic aspects such as intentional and contextual factors. Looking ahead, we envision extending this framework to new modalities and domains and implementing it in computational experiments.